OpenStudy (anonymous):

find the arithmetic means in the given sequence. 175, ?, ?, ?, 235.

OpenStudy (anonymous):

205

OpenStudy (amistre64):

i would shift the range by subtracting 175 from all of it 0, ?, ?, ?, 60 now we need to determine 3 numbers that are evenly spaced within this range

OpenStudy (amistre64):

60/2 = 30 for the middle and 30/2 = 15 for the first one 30+15 = 45 for the third one 0,15,30,45,60 now add back the 175

OpenStudy (amistre64):

i spose 60/(3+1) would have been sufficient for a common difference